The length of a rectangle is 3 feet greater than its width. The area of the rectangle is 108ft square. Find the length and width.
Answer by jorel1380(3719)

x(x+3)=108
x2+3x=108
x2+3x-108=0
(x+12)(x-9)=0
x=-12 or 9
The length and width is 9'x12'..